OpenStudy (anonymous):

A grocery store sells pre-cut fruit. Bowls of one type of fruit are four dollars. Bowls of assorted fruit are six dollars. In one month, Benjamin paid eighty-four dollars for 16 bowls of fruit. How many assorted fruit bowls did he buy? How do you set the equations up for this? A grocery store sells pre-cut fruit. Bowls of one type of fruit are four dollars. Bowls of assorted fruit are six dollars. In one month, Benjamin paid eighty-four dollars for 16 bowls of fruit. How many assorted fruit bowls did he buy? How do you set the equations up for this? @Mathematics

OpenStudy (anonymous):

first time doing these

OpenStudy (anonymous):

dont need an answer just how to set it up and work it. Please

OpenStudy (anonymous):

I think 10

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Bowl's of one type of fruit = x Bowl's of assorted fruit = y x+y=16 x = 16-y 4x+6y = 84 4(16-y) + 6y = 84 64-4y+6y = 84 64+2y = 84 2y = 84-64 2y = 20 y = 10